find the zeros of each function;
y=x(x+5)(x-8)

Found 2 solutions by stanbon, ReadingBoosters:
Answer by stanbon(75887)

find the zeros of each function;
y=x(x+5)(x-8)
-----
Solve x(x+5)(x-8) = 0
x = 0 or x = -5 or x = 8
============================
Cheers,
Stan H.
===========

Answer by ReadingBoosters(3246)

0= x(x+5)(x-8)
...
x = 0
x + 5 = 0; so x = -5
x - 8 = 0; so x = 8
...
roots at -5, 0 and 8
..............................
Delighted to help; feel free to email directly at HomeworkHelpers@ReadingBoosters.com.
-Reading Boosters
Website: www.MyHomeworkAnswers.com